CH_{2} ANTISYMMETRIC STRETCHING VIBRATIONS OF THE ALLYL RADICAL

DAIRENE UY, \underline{SCOTT DAVIS}, DAVID J. NESBITT, JILA, National Institute of Standards and Technology and Department of Chemistry and Biochemistry, University of Colorado, Boulder, Colorado 80309-0440.

A glow discharge, slit supersonic expansion in conjunction with direct infrared laser absorption have been utilized to record high resolution vibration-rotation spectra of the C_{3}H_{5} allyl radical. The slit supersonic expansion results in efficient rotational cooling (T_{rot} ≤ 20K),thereby facilitating assignment. Specifically, approximately 50 transitions are assigned for both the in-phase (ν_{1}) and out-of-phase (ν_{13}) CH_{2} antisymmetric vibrations. Least squares fits of the transition frequencies to an asymmetric top Hamiltonian provide both ground and excited state rotational constants. While the overall quality of the fits (σ ≈ 5 × 10^{−4} cm^{−1}) are good, residuals do indicate perturbations in the vibrationally excited state. Due to the high instrumental resolution (slit supersonic Doppler width ≈ 70 MHz) spin-rotation broadening is observed in several low J transitions. Small step size (4 MHz) scans over selected transitions coupled with a detailed lineshape analysis indicate a spin-rotation constant ε_{aa} of -47(4) MHz.